In a study of 795 randomly selected medical malpractice lawsuits, it was found that 481 of them were dropped or dismissed. Use a 0.05 significance level to test the claim that most medical malpractice lawsuits are dropped or dismissed.

Which of the following is the hypothesis test to be conducted?

A.

Upper H 0 : p equals 0.5H0: p=0.5

Upper H 1 : p greater than 0.5H1: p>0.5

B.

Upper H 0 : p less than 0.5H0: p<0.5

Upper H 1 : p equals 0.5H1: p=0.5

C.

Upper H 0 : p equals 0.5H0: p=0.5

Upper H 1 : p not equals 0.5H1: p≠0.5

D.

Upper H 0 : p equals 0.5H0: p=0.5

Upper H 1 : p less than 0.5H1: p<0.5

E.

Upper H 0 : p not equals 0.5H0: p≠0.5

Upper H 1 : p equals 0.5H1: p=0.5

F.

Upper H 0 : p greater than 0.5H0: p>0.5

Upper H 1 : p equals 0.5H1: p=0.5

What is the test statistic?

A.

5.92

B.

-5.92

C.

6.06

D.

correct answer is not given

What is the critical value?

A.

1.645

B.

-1.645

C.

correct answer is not given

D.

1.96

What is the conclusion about the null hypothesis?

A.

Fail to reject the null hypothesis

B.

Accept the null hypothesis

C.

Reject the null hypothesis

What is the final conclusion?

A.

There is not sufficient evidence to warrant rejection of the claim that most medical malpractice lawsuits are dropped or dismissed.

B.

There is sufficient evidence to support the claim that most medical malpractice lawsuits are dropped or dismissed.

C.

There is sufficient evidence to warrant rejection of the claim that most medical malpractice lawsuits are dropped or dismissed.

D.

There is not sufficient evidence to support the claim that most medical malpractice lawsuits are dropped or dismissed.
Answer #1

Here' the answer to the question. Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

Which of the following is the hypothesis test to be conducted?

p' = x/n = 481/795 = 0.605

We have to test the claim that most medical malpractice lawsuits are dropped or dismissed.

So, our altenrative hypothesis becomes the claim: Ha: p>.50

Hence, Ho: p=.50


What is the test statistic?

A. is correct
test statistic = (p-p0)/sqrt(p0*p0'/n) = (.605-.50)/sqrt(.5*.5/795) = 5.92

A. is correct. It' 5.92

What is the critical value?

Critical value is basically at the Z value for 95% confidence interval , which is +1.96
(since its a right tailed test)

What is the conclusion about the null hypothesis?

We C. reject the null hypothesis ( since our test statistic is more than critical value of test statistic)

What is the final conclusion?

The conclusion is : Since our test-statistic is more than t-critical value(5.92>1.96), we
say that B. There is sufficient evidence to support the claim that most medical malpractice lawsuits are dropped or dismissed.